# Dataset Distillation via Knowledge Distillation: Towards Efficient Self-Supervised Pre-training of Deep Networks

### v1

@@ 4.2 Matching Knowledge Distillation Trajectories
To reduce the length and variance of SSL trajectories, our key idea is to leverage the recent results in knowledge distillation (KD) [17]. We first introduce KD, and then discuss our method, MKDT, Matching Knowledge Distillation Trajectories, that leverages KD to reduce the length and variance of SSL trajectories.
Knowledge Distillation (KD) KD refers to distilling the knowledge of a larger model (teacher) into a smaller model (student) to achieve similar generalization as the larger model, but with reduced complexity and faster inference times. Here, we rely on the knowledge distillation objective for SSL models, introduced in [41]:
$\mathcal{L_{KD}}=\mathbb{E}_{x\sim\mathcal{D}_{\mathsf{real}}}\big{[}\text{MSE}(f_{S}(x),f_{T}(x))\big{]},\vspace{-4mm}$ (5)
where $f_{S}$ and $f_{T}$ represent the student and teacher encoders respectively. [41] trains student models with the aforementioned KD objective and the original SSL Loss. However, we only minimize the MSE between student and teacher representations to avoid the issues with matching SSL training trajectories (discussed in Sec. 4.1).
@@
In our experiments, $\mathcal{L}_{\mathsf{SSL}}$ is the BarlowTwins loss function shown in equation 1, but our method is agnostic to the choice of SSL algorithm. Since KD with larger models leads to better downstream performance [13], we use a teacher model that is much larger than the student encoder used for creating the expert trajectories for distillation. For example, in our experiments we use a ResNet-18 as the teacher encoder and a 3 or 4-layer ConvNet as the student encoder.
@@ 5 Experiments
CIFAR100 1% Random Subset $42.45\scriptstyle{\pm 0.76}$ $11.55\scriptstyle{\pm 0.37}$ $4.17\scriptstyle{\pm 0.15}$ $2.47\scriptstyle{\pm 0.17}$ $1.51\scriptstyle{\pm 0.12}$ $2.12\scriptstyle{\pm 0.18}$ $\textbf{2.61}\scriptstyle{\pm 0.61}$
@@
Random $18.80\scriptstyle{\pm 0.58}$ $44.24\scriptstyle{\pm 0.85}$ $10.33\scriptstyle{\pm 0.16}$ $2.15\scriptstyle{\pm 0.34}$ $1.81\scriptstyle{\pm 0.16}$ $2.52\scriptstyle{\pm 0.29}$ $5.41\scriptstyle{\pm 0.61}$
MKDT $\textbf{23.23}\scriptstyle{\pm 0.79}$ $\textbf{49.13}\scriptstyle{\pm 0.69}$ $\textbf{13.35}\scriptstyle{\pm 0.24}$ $\textbf{1.68}\scriptstyle{\pm 0.10}$ $\textbf{1.67}\scriptstyle{\pm 0.09}$ $\textbf{2.64}\scriptstyle{\pm 0.16}$ $\textbf{6.15}\scriptstyle{\pm 0.47}$
@@
MKDT $\textbf{21.51}\scriptstyle{\pm 0.17}$ $\textbf{46.10}\scriptstyle{\pm 0.60}$ $\textbf{11.57}\scriptstyle{\pm 0.17}$ $\textbf{2.05}\scriptstyle{\pm 0.43}$ $\textbf{1.86}\scriptstyle{\pm 0.05}$ $\textbf{2.36}\scriptstyle{\pm 0.29}$ $\textbf{5.17}\scriptstyle{\pm 0.93}$
@@
CIFAR10 1% Random $35.20\scriptstyle{\pm 1.12}$ $7.35\scriptstyle{\pm 0.28}$ $2.29\scriptstyle{\pm 0.14}$ $2.21\scriptstyle{\pm 0.09}$ $1.19\scriptstyle{\pm 0.06}$ $1.83\scriptstyle{\pm 0.16}$ $1.83\scriptstyle{\pm 0.24}$
MKDT $\textbf{40.77}\scriptstyle{\pm 1.05}$ $\textbf{9.17}\scriptstyle{\pm 0.13}$ $\textbf{3.06}\scriptstyle{\pm 0.16}$ $\textbf{2.69}\scriptstyle{\pm 0.21}$ $\textbf{1.35}\scriptstyle{\pm 0.06}$ $\textbf{2.02}\scriptstyle{\pm 0.23}$ $\textbf{1.88}\scriptstyle{\pm 0.22}$
@@
MKDT $\textbf{51.77}\scriptstyle{\pm 0.25}$ $\textbf{18.07}\scriptstyle{\pm 0.52}$ $\textbf{6.55}\scriptstyle{\pm 0.23}$ $\textbf{5.90}\scriptstyle{\pm 0.76}$ $\textbf{1.89}\scriptstyle{\pm 0.15}$ $\textbf{2.98}\scriptstyle{\pm 0.36}$ $\textbf{4.09}\scriptstyle{\pm 0.28}$
CIFAR100 1% Random $34.67\scriptstyle{\pm 0.89}$ $7.35\scriptstyle{\pm 0.54}$ $2.29\scriptstyle{\pm 0.16}$ $2.23\scriptstyle{\pm 0.21}$ $1.10\scriptstyle{\pm 0.06}$ $1.78\scriptstyle{\pm 0.26}$ $1.77\scriptstyle{\pm 0.13}$
@@
In this section, we evaluate the downstream generalization of models trained using the synthetic sets distilled by MKDT that are 2% and 5% of the original dataset’s size for CIFAR10, CIFAR100 [19], TinyImageNet [20]. We also conduct ablation studies over initialization of the distilled set and the SSL algorithm. Finally, we also consider the generalization of the distilled sets to larger architectures.
Distillation Setup We use ResNet-18 trained with Barlow Twins [43] as the teacher encoder and train $K=100$ student encoders (using ConvNets) to generate the expert trajectories. As in previous work [1; 44; 46; 3; 9; 6], we use a 3-layer ConvNet for the lower resolution CIFAR datasets and a 4-layer ConvNet for the higher resolution TinyImageNet. Exact hyperparamters in Appendix A.
@@
Baselines We compare pre-training with MKDT distilled sets to pre-training with random subsets, SAS subsets [14], KRR-ST distilled sets [21] as well as pre-training on the full data. For KRR-ST, we use the provided distilled sets for CIFAR100 and Tiny ImageNet, and distill using the provided code for CIFAR10. We omit SAS for TinyImageNet as this subset was not provided. For distilled sets of size 5%, we consider only Random and MKDT , since other baselines did not provide distilled sets.
@@
Downstream Generalization Performance Table 1 demonstrates that pre-training on CIFAR10 using MKDT with a 2% distilled set improves performance by 8% on CIFAR10 and 5% on downstream tasks over the KRR-ST baseline. Gains on CIFAR10 are consistent across 1% and 5% labeled data, but improvements on downstream tasks are more pronounced with 5%, indicating MKDT scales well with more labeled data. On CIFAR100, MKDT 2% distilled set improves performance by 6% and 8% on CIFAR100 and downstream tasks, respectively. Additionally, MKDT shows up to 3% improvement on downstream tasks for larger, higher-resolution datasets like TinyImageNet (200K examples, 64x64 resolution), highlighting MKDT’s scalability. KRR-ST consistently fails to outperform SSL pre-training on random subsets across all settings. In Appendix B, we verify that this holds for fine-tuning experiments from KRR-ST [21], affirming MKDT as the only effective DD method for SSL pre-training. Table 4 shows that pre-training with larger distilled sets (5% of full data) further enhances performance by up to 13%, confirming MKDT scales effectively with distilled set size as well. Table 7 shows that MKDT outperforms the strongest baseline (random subsets) by 5% on pre-training and 7% on downstream tasks when using 10% and 50% labeled data.
@@
Ablations We perform ablations over two factors: 1) initialization and 2) SSL algorithm. Table 2 presents results for pre-training with MKDT using random subset initialization, as well as results for pre-training directly on the high-loss subset initialization used by MKDT. Interestingly, the KD objective for SSL pre-training benefits slightly more from the high-loss subset than from random subsets. Consequently, MKDT initialized from the high-loss subset performs better than when initialized from a random subset. Table 6 shows results for MKDT using a teacher model trained with SimCLR [2] instead of BarlowTwins. Specifically, we train a ResNet-18 for 400 epochs using SimCLR. Here too, MKDT achieves approximately 6% higher performance compared to random subsets across downstream datasets. This confirms that MKDT generalizes across different SSL training algorithms.
@@
Generalization to Larger Architectures Table 5 compares CIFAR100 5% size MKDT distilled set to 5% size random subsets, using the larger ResNet-10 and ResNet-18 architectures. Across all downstream tasks, we confirm that MKDT distilled sets outperform baselines even when using larger architectures. Surprisingly, the larger ResNet-18 slightly under-performs the smaller ResNet-10. This trend is observed across all baselines, including no-pretraining. We conjecture this is due to larger models requiring a lot more data to be able to use their extra capacity to surpass their smaller counterparts.

### v3

@@ 4.2 Matching Knowledge Distillation Trajectories
To reduce the length and variance of SSL trajectories, our key idea is to leverage the recent results in knowledge distillation (KD) Kim et al. (2021). We first introduce KD, and then discuss our method, MKDT, Matching Knowledge Distillation Trajectories, that leverages KD to reduce the length and variance of SSL trajectories.
Knowledge Distillation (KD) KD refers to distilling the knowledge of a larger model (teacher) into a smaller model (student) to achieve similar generalization as the larger model, but with reduced complexity and faster inference times. Here, we rely on the knowledge distillation objective for SSL models, introduced in Yu et al. (2019):
$\mathcal{L_{KD}}=\mathbb{E}_{x\sim\mathcal{D}_{\mathsf{real}}}\big{[}\text{MSE}(f_{S}(x),f_{T}(x))\big{]},$ (5)
where $f_{S}$ and $f_{T}$ represent the student and teacher encoders respectively. (Yu et al., 2019) trains student models with the aforementioned KD objective and the original SSL Loss. However, we only minimize the MSE between student and teacher representations to avoid the issues with matching SSL training trajectories (discussed in Sec. 4.1).
@@
In our experiments, $\mathcal{L}_{\mathsf{SSL}}$ is the BarlowTwins loss function shown in equation 1, but our method is agnostic to the choice of SSL algorithm. Since KD with larger models leads to better downstream performance (Huang et al., 2022), we use a teacher model that is much larger than the student encoder used for creating the expert trajectories for distillation. For example, in our experiments we use a ResNet-18 as the teacher encoder and a 3 or 4-layer ConvNet as the student encoder.
@@ 5 Experiments
KRR-ST ${36.69\scriptstyle{\pm 0.88}}$ ${8.69\scriptstyle{\pm 0.32}}$ ${3.20\scriptstyle{\pm 0.23}}$ ${2.26\scriptstyle{\pm 0.13}}$ ${1.33\scriptstyle{\pm 0.09}}$ ${1.91\scriptstyle{\pm 0.34}}$ ${2.39\scriptstyle{\pm 0.18}}$
@@
KRR-ST ${47.40\scriptstyle{\pm 0.34}}$ ${16.95\scriptstyle{\pm 0.53}}$ ${7.10\scriptstyle{\pm 0.27}}$ ${5.56\scriptstyle{\pm 0.77}}$ ${1.98\scriptstyle{\pm 0.07}}$ ${2.78\scriptstyle{\pm 0.16}}$ ${4.38\scriptstyle{\pm 0.04}}$
@@
CIFAR100 1% Random Subset $42.45\scriptstyle{\pm 0.76}$ $11.55\scriptstyle{\pm 0.37}$ $4.17\scriptstyle{\pm 0.15}$ $2.47\scriptstyle{\pm 0.17}$ $1.51\scriptstyle{\pm 0.12}$ $2.12\scriptstyle{\pm 0.18}$ $2.61\scriptstyle{\pm 0.61}$
KRR-ST ${37.86\scriptstyle{\pm 1.14}}$ ${9.02\scriptstyle{\pm 0.24}}$ ${2.94\scriptstyle{\pm 0.13}}$ ${2.42\scriptstyle{\pm 0.35}}$ ${1.50\scriptstyle{\pm 0.07}}$ ${1.99\scriptstyle{\pm 0.19}}$ ${\textbf{3.04}\scriptstyle{\pm 0.36}}$
@@
KRR-ST ${47.53\scriptstyle{\pm 0.11}}$ ${17.24\scriptstyle{\pm 0.47}}$ ${6.60\scriptstyle{\pm 0.32}}$ ${5.37\scriptstyle{\pm 0.85}}$ ${2.31\scriptstyle{\pm 0.33}}$ ${2.87\scriptstyle{\pm 0.27}}$ ${5.23\scriptstyle{\pm 0.14}}$
@@
KRR-ST ${34.27\scriptstyle{\pm 1.36}}$ ${7.54\scriptstyle{\pm 0.35}}$ ${3.19\scriptstyle{\pm 0.22}}$ ${2.11\scriptstyle{\pm 0.23}}$ ${1.30\scriptstyle{\pm 0.12}}$ ${1.68\scriptstyle{\pm 0.20}}$ ${2.65\scriptstyle{\pm 0.64}}$
@@
KRR-ST ${42.82\scriptstyle{\pm 0.46}}$ ${13.71\scriptstyle{\pm 0.30}}$ ${6.50\scriptstyle{\pm 0.23}}$ ${4.36\scriptstyle{\pm 0.49}}$ ${1.97\scriptstyle{\pm 0.06}}$ ${2.75\scriptstyle{\pm 0.37}}$ ${3.97\scriptstyle{\pm 0.14}}$
@@
Random $18.80\scriptstyle{\pm 0.58}$ $44.24\scriptstyle{\pm 0.85}$ $10.33\scriptstyle{\pm 0.16}$ $\textbf{2.15}\scriptstyle{\pm 0.34}$ $\textbf{1.81}\scriptstyle{\pm 0.16}$ $2.52\scriptstyle{\pm 0.29}$ $5.41\scriptstyle{\pm 0.61}$
KRR-ST ${13.84\scriptstyle{\pm 0.78}}$ ${39.21\scriptstyle{\pm 0.55}}$ ${8.04\scriptstyle{\pm 0.52}}$ ${2.12\scriptstyle{\pm 0.15}}$ ${1.16\scriptstyle{\pm 0.05}}$ ${1.77\scriptstyle{\pm 0.14}}$ ${4.56\scriptstyle{\pm 0.42}}$
MKDT $\textbf{23.23}\scriptstyle{\pm 0.79}$ $\textbf{49.13}\scriptstyle{\pm 0.69}$ $\textbf{13.35}\scriptstyle{\pm 0.24}$ $1.68\scriptstyle{\pm 0.10}$ $1.67\scriptstyle{\pm 0.09}$ $\textbf{2.64}\scriptstyle{\pm 0.16}$ $\textbf{6.15}\scriptstyle{\pm 0.47}$
@@
KRR-ST ${12.30\scriptstyle{\pm 0.83}}$ ${35.73\scriptstyle{\pm 1.07}}$ ${7.21\scriptstyle{\pm 0.35}}$ ${\textbf{2.32}\scriptstyle{\pm 0.39}}$ ${1.18\scriptstyle{\pm 0.16}}$ ${1.81\scriptstyle{\pm 0.14}}$ ${2.45\scriptstyle{\pm 0.12}}$
MKDT $\textbf{21.51}\scriptstyle{\pm 0.17}$ $\textbf{46.10}\scriptstyle{\pm 0.60}$ $\textbf{11.57}\scriptstyle{\pm 0.17}$ $2.05\scriptstyle{\pm 0.43}$ $\textbf{1.86}\scriptstyle{\pm 0.05}$ $\textbf{2.36}\scriptstyle{\pm 0.29}$ $\textbf{5.17}\scriptstyle{\pm 0.93}$
@@
CIFAR10 1% Random Subset $35.20\scriptstyle{\pm 1.12}$ $7.35\scriptstyle{\pm 0.28}$ $2.29\scriptstyle{\pm 0.14}$ $2.21\scriptstyle{\pm 0.09}$ $1.19\scriptstyle{\pm 0.06}$ $1.83\scriptstyle{\pm 0.16}$ $1.83\scriptstyle{\pm 0.24}$
KRR-ST ${36.90\scriptstyle{\pm 1.30}}$ ${8.38\scriptstyle{\pm 0.17}}$ ${2.95\scriptstyle{\pm 0.12}}$ ${2.45\scriptstyle{\pm 0.13}}$ ${1.19\scriptstyle{\pm 0.09}}$ ${1.87\scriptstyle{\pm 0.18}}$ ${\textbf{2.35}\scriptstyle{\pm 0.06}}$
MKDT $\textbf{40.77}\scriptstyle{\pm 1.05}$ $\textbf{9.17}\scriptstyle{\pm 0.13}$ $\textbf{3.06}\scriptstyle{\pm 0.16}$ $\textbf{2.69}\scriptstyle{\pm 0.21}$ $\textbf{1.35}\scriptstyle{\pm 0.06}$ $\textbf{2.02}\scriptstyle{\pm 0.23}$ $1.88\scriptstyle{\pm 0.22}$
@@
KRR-ST ${46.87\scriptstyle{\pm 0.52}}$ ${16.29\scriptstyle{\pm 0.37}}$ ${6.31\scriptstyle{\pm 0.43}}$ ${5.31\scriptstyle{\pm 0.63}}$ ${\textbf{1.89}\scriptstyle{\pm 0.14}}$ ${2.66\scriptstyle{\pm 0.18}}$ ${\textbf{4.36}\scriptstyle{\pm 0.16}}$
MKDT $\textbf{51.77}\scriptstyle{\pm 0.25}$ $\textbf{18.07}\scriptstyle{\pm 0.52}$ $\textbf{6.55}\scriptstyle{\pm 0.23}$ $\textbf{5.90}\scriptstyle{\pm 0.76}$ $\textbf{1.89}\scriptstyle{\pm 0.15}$ $\textbf{2.98}\scriptstyle{\pm 0.36}$ $4.09\scriptstyle{\pm 0.28}$
CIFAR100 1% Random Subset $34.67\scriptstyle{\pm 0.89}$ $7.35\scriptstyle{\pm 0.54}$ $2.29\scriptstyle{\pm 0.16}$ $2.23\scriptstyle{\pm 0.21}$ $1.10\scriptstyle{\pm 0.06}$ $1.78\scriptstyle{\pm 0.26}$ $1.77\scriptstyle{\pm 0.13}$
KRR-ST ${36.57\scriptstyle{\pm 1.02}}$ ${8.38\scriptstyle{\pm 0.36}}$ ${3.01\scriptstyle{\pm 0.22}}$ ${2.41\scriptstyle{\pm 0.15}}$ ${1.28\scriptstyle{\pm 0.02}}$ ${1.71\scriptstyle{\pm 0.30}}$ ${1.98\scriptstyle{\pm 0.24}}$
@@
KRR-ST ${46.76\scriptstyle{\pm 0.50}}$ ${15.75\scriptstyle{\pm 0.46}}$ ${6.17\scriptstyle{\pm 0.20}}$ ${5.43\scriptstyle{\pm 0.65}}$ ${\textbf{1.93}\scriptstyle{\pm 0.06}}$ ${2.61\scriptstyle{\pm 0.25}}$ ${3.55\scriptstyle{\pm 0.29}}$
@@
KRR-ST ${51.02\pm\scriptstyle{0.53}}$ ${21.01\pm\scriptstyle{0.14}}$ ${8.95\pm\scriptstyle{0.26}}$ ${7.91\pm\scriptstyle{0.76}}$ ${2.44\pm\scriptstyle{0.12}}$ ${3.54\pm\scriptstyle{0.29}}$ ${6.82\pm\scriptstyle{0.64}}$
@@
KRR-ST ${58.09\pm\scriptstyle{0.07}}$ ${29.01\pm\scriptstyle{0.30}}$ ${15.94\pm\scriptstyle{0.31}}$ ${17.60\pm\scriptstyle{1.04}}$ ${5.01\pm\scriptstyle{0.34}}$ ${6.81\pm\scriptstyle{0.35}}$ ${15.92\pm\scriptstyle{0.64}}$
@@
KRR-ST ${52.40\pm\scriptstyle{0.73}}$ ${21.39\pm\scriptstyle{0.16}}$ ${8.21\pm\scriptstyle{0.07}}$ ${7.64\pm\scriptstyle{0.36}}$ ${2.34\pm\scriptstyle{0.12}}$ ${3.76\pm\scriptstyle{0.25}}$ ${6.52\pm\scriptstyle{1.28}}$
@@
KRR-ST ${58.57\pm\scriptstyle{0.78}}$ ${29.46\pm\scriptstyle{0.81}}$ ${15.70\pm\scriptstyle{0.07}}$ ${15.89\pm\scriptstyle{0.54}}$ ${4.82\pm\scriptstyle{0.41}}$ ${7.00\pm\scriptstyle{0.18}}$ ${15.07\pm\scriptstyle{0.46}}$
@@
KRR-ST ${45.48\pm\scriptstyle{0.84}}$ ${17.02\pm\scriptstyle{0.26}}$ ${8.88\pm\scriptstyle{0.41}}$ ${5.29\pm\scriptstyle{0.16}}$ ${2.08\pm\scriptstyle{0.21}}$ ${3.21\pm\scriptstyle{0.15}}$ ${6.10\pm\scriptstyle{1.44}}$
@@
KRR-ST ${48.16\pm\scriptstyle{1.18}}$ ${20.01\pm\scriptstyle{0.40}}$ ${13.59\pm\scriptstyle{0.47}}$ ${8.66\pm\scriptstyle{1.29}}$ ${3.46\pm\scriptstyle{0.22}}$ ${4.98\pm\scriptstyle{0.31}}$ ${15.12\pm\scriptstyle{0.37}}$
@@
In this section, we evaluate the downstream generalization of models trained using the synthetic sets distilled by MKDT that are 2% and 5% of the original dataset’s size for CIFAR10, CIFAR100 (Krizhevsky & Hinton, 2009), TinyImageNet (Le & Yang, 2015). We also conduct ablation studies over initialization of the distilled set and the SSL algorithm. Finally, we also consider the generalization of the distilled sets to larger architectures.
Distillation Setup We use ResNet-18 trained with Barlow Twins (Zbontar et al., 2021) as the teacher encoder and train $K=100$ student encoders (using ConvNets) to generate the expert trajectories. As in previous work (Cazenavette et al., 2022; Zhao & Bilen, 2021; Zhao et al., 2021; Chen et al., 2023; Du et al., 2023; Cui et al., 2022), we use a 3-layer ConvNet for the lower resolution CIFAR datasets and a 4-layer ConvNet for the higher resolution TinyImageNet. Exact hyperparamters in Appendix A.
@@
Baselines We compare pre-training with MKDT distilled sets to pre-training with random subsets, SAS subsets (Joshi & Mirzasoleiman, 2023), KRR-ST distilled sets (Lee et al., 2023) as well as pre-training on the full data. For KRR-ST, we use the provided distilled sets for CIFAR100 and Tiny ImageNet, and distill using the provided code for CIFAR10. We omit SAS for TinyImageNet as this subset was not provided. For distilled sets of size 5%, we consider only Random and MKDT, since other baselines did not provide distilled sets.
@@
Downstream Generalization Performance Table 1 demonstrates that pre-training on CIFAR10 using MKDT with a 2% distilled set improves performance by 8% on CIFAR10 and 5% on downstream tasks over the KRR-ST baseline. Gains on CIFAR10 are consistent across 1% and 5% labeled data, but improvements on downstream tasks are more pronounced with 5%, indicating MKDT scales well with more labeled data. On CIFAR100, MKDT 2% distilled set improves performance by 6% and 8% on CIFAR100 and downstream tasks, respectively. Additionally, MKDT shows up to 3% improvement on downstream tasks for larger, higher-resolution datasets like TinyImageNet (200K examples, 64x64 resolution), highlighting MKDT’s scalability. KRR-ST consistently fails to outperform SSL pre-training on random subsets across all settings. In Appendix B, we verify that this holds for fine-tuning experiments from KRR-ST (Lee et al., 2023), affirming MKDT as the only effective DD method for SSL pre-training. Table 4 shows that pre-training with larger distilled sets (5% of full data) further enhances performance by up to 13%, confirming MKDT scales effectively with distilled set size as well. Table 7 shows that MKDT outperforms the strongest baseline (random subsets) by 5% on pre-training and 7% on downstream tasks when using 10% and 50% labeled data.
@@
Ablations We perform ablations over two factors: 1) initialization and 2) SSL algorithm. Table 2 presents results for pre-training with MKDT using random subset initialization, as well as results for pre-training directly on the high-loss subset initialization used by MKDT. Interestingly, the KD objective for SSL pre-training benefits slightly more from the high-loss subset than from random subsets. Consequently, MKDT initialized from the high-loss subset performs better than when initialized from a random subset. Table 6 shows results for MKDT using a teacher model trained with SimCLR (Chen et al., 2020) instead of BarlowTwins. Specifically, we train a ResNet-18 for 400 epochs using SimCLR. Here too, MKDT achieves approximately 6% higher performance compared to random subsets across downstream datasets. This confirms that MKDT generalizes across different SSL training algorithms.
@@
Generalization to Larger Architectures Table 5 compares CIFAR100 5% size MKDT distilled set to 5% size random subsets, using the larger ResNet-10 and ResNet-18 architectures. Across all downstream tasks, we confirm that MKDT distilled sets outperform baselines even when using larger architectures. Surprisingly, the larger ResNet-18 slightly under-performs the smaller ResNet-10. This trend is observed across all baselines, including no pre-training. We conjecture this is due to larger models requiring a lot more data to be able to use their extra capacity to surpass their smaller counterparts.